OpenStudy (anonymous):

Which percent is equal to the given fraction? 1/20. A. 0.5% B. 5% C. 10% D. 20%

pooja195 (pooja195):

Just divide 1/20=?

OpenStudy (anonymous):

0.05?

OpenStudy (alex6799):

5%

pooja195 (pooja195):

Right you would move the decimal to the end to get your percentage (move to the left till you have a whole number)

pooja195 (pooja195):

Welcome to openstudy :-)

OpenStudy (bibby):

percent means parts out of 100, set up a proportion: \(\large \dfrac{1}{20}=\dfrac{x}{100}\)

OpenStudy (anonymous):

thank you :3 i'll medal if we were right

OpenStudy (kaaaykillaaah):

For future students looking for a direct answer it s A